OpenStudy (anonymous):

A monument is built out of toothpicks. There are 30 rows and each row has seven more toothpicks than the row above. The top row has five toothpicks. How many toothpicks are on the bottom row? Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (lgbasallote):

procedure taken -> set-up toothpicks

OpenStudy (zepp):

I would be an arithmetic series, from top, you'll have 5, 5+7, 5+2*7+5+3*7 etc etc

OpenStudy (zepp):

It*

OpenStudy (anonymous):

yea but I got 208, is that right?

OpenStudy (zepp):

To find any term in an arithmetic series, you'll use the formula \(a_{n} = a_{1}+(n+1)d\) Since we do not have to first term, we'll work backward. \(\huge 5 = a_{1}+(30+1)(-7)\) We have here \(a_{30}\) = 5, 30th term, common difference being -7. \(\huge 5 = a_{1}+217\) \(\huge a_{1} = 222\)

OpenStudy (anonymous):

but u have a a1, its 5

OpenStudy (anonymous):

a1 = 5, n = 30, and d = 7, right?

OpenStudy (zepp):

Nuh-uh, 5 is the last term, you have 30 row and 5 is the top row mean that ...,...,...,...,...,...,...,...,...,..., (29 terms later), 5

OpenStudy (zepp):

You can work that way, too, I think.

OpenStudy (anonymous):

but it says top row, which means the first row, right?

OpenStudy (zepp):

Uhhh, you are right :O

OpenStudy (anonymous):

really? s its 208?

OpenStudy (zepp):

So yes top row, first term = 5 Common difference +7 We are looking for the 30th term a30 = 5+(30+1)7 a30 = 222 If would still be 222 :P

OpenStudy (anonymous):

its minus, the formula is an = a1 + (n - 1)d and the the 30 is what we are trying to find because of the last, 30th row... right?

OpenStudy (zepp):

Right, oh yeah, the minus >.> So yes, your answer would be 208 (omg so misleading people today lol)

OpenStudy (anonymous):

haha its okay, thank u so much for talking it through with me! :)

OpenStudy (zepp):

Haha, no problem :)